OpenStudy (anonymous):

how do I find the inverse of the function f(x)= 4x^3 -3 algebraically?

OpenStudy (mathstudent55):

1. Replace f(x) with y. 2. Switch x and y. 3. Solve for y. 4. Replace y with \(f^{-1}(x) \).

OpenStudy (muzzack):

try to solve for x Once you have x , you'll switch x and y; the y is the inverse.

OpenStudy (anonymous):

what about the x^3?

geerky42 (geerky42):

Same steps explained above. for \(x^3\), you have \[y=x^3\\ x = y^3\\ \sqrt[3]{x}=y \\ \boxed{\sqrt[3]{x}=f^{-1}(x)}\]

OpenStudy (anonymous):

y=4x^3-3 x=4y^3-3 x+3=4y^3 x+3/4=y^3 the cube root\[\sqrt[3]{?x+3/4}=y\]

OpenStudy (anonymous):

the whole cube root is over 4 tho not just the 3 which is how it looks cuz I tried to use the equations thingie

OpenStudy (mathstudent55):

\( f(x)= 4x^3 -3\) 1. \( y = 4x^3 -3\) 2. \( x = 4y^3 -3\) 3. \( x = 4y^3 -3\) \(x + 3= 4y^3 \) \(4y^3 = x + 3 \) \(y^3 = \dfrac{x + 3}{4} \) \(y = \sqrt[3]{ \dfrac{x + 3}{4} } \) 4. \(f^{-1}(x) = \sqrt[3]{ \dfrac{x + 3}{4} } \)